### v2

# Non-equilibrium dynamics of localization phase transition in the non-Hermitian Disorder-Aubry-André model

Abstract
The driven dynamics of localization transitions in a non-Hermitian Disorder-Aubry-André (DAA) model are thoroughly examined under both open boundary conditions (OBC) and periodic boundary conditions (PBC). Through an analysis of the static properties of observables, including the localization length ($\xi$), inverse participation ratio ($\rm IPR$), and energy gap ($\Delta E$), we found that the critical exponents examined under PBC are also applicable under OBC. The Kibble-Zurek scaling (KZS) for the driven dynamics in the non-Hermitian DAA systems is formulated and numerically verified for different quench directions. Notably, for the dynamical paths considered, boundary conditions had minimal impact on the evolution process. This study generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.
I Introduction
The study of localization phase transition behaviors is a longstanding research topic in condensed matter physics Anderson1958 ; Aubry1980 ; Biddle070601 ; Xu2019 ; Khemani2017 ; HepengYao2019 ; Schirmann2024 ; Ribeiro2024 ; Jan2023 ; Alexey2023 ; Agrawal2020 ; Agrawalprl2020 ; Goblot2020 ; Roy2022 ; Agrawal2022 ; Zhou2023 ; Smith176601 . Besides the Anderson model Anderson1958 , the localization phase transition has also been found in the quasi-periodic systems, and various quasi-periodic models have been proposed, such as the Fibonacci model Strkalj2021 ; zhai2021 ; Guo2024 , the Aubry-André (AA) model and its various extensions Aubry1980 ; Sarma1988 ; Biddle2011 ; Wangprl2020 ; Liu2021 ; Gao2024 ; Ting2022 ; Longhi20192 ; Zeng2024 ; Wang2021 ; Borgnia1 ; GYSun2022 . Studies have shown that there are certain differences in the localization phase transition between Anderson model and quasi-periodic models. For example, the localization phase transition can even occur for the one-dimensional (1D) AA model Aubry1980 ; Jan2023 ; Alexey2023 , and the critical exponents for the Anderson and AA models belong to two different universality classes Agrawal2020 ; Wei2019 ; Wang014206 ; Cestari2011 ; Xuan2022 . On the other hand, significant advancements in simulating nonequilibrium dynamics in quantum systems have been made Wangxh2024 ; ChenJJ2024 ; Cheng2399 ; Xue2024 ; Zhang2023131 ; You2024109 ; Mak2024 ; Lizhen2024 . Considerable attention has also been given to the nonequilibrium dynamics in localization transitions Yang2017 ; Heyl2018 ; Bertini2024 ; Khan2024 . Several protocols have been employed to drive a quasi-periodic or disorder system out of equilibrium, such as suddenly changing the Hamiltonian Yang2017 ; Yin2018 ; Modak2021 , introducing a time-(quasi)periodic driving Yang2019 ; Zhou2021 ; ZhouDu2021 , or linearly quenching across the localization transition point Sinha2019 ; Liang2024 ; Xuan2023 . Sudden quenching of the Anderson and AA models, with initial and post-quench Hamiltonians in different phases, will result in periodic zeros in the Loschmidt echo, indicating the localization-delocalization transition Yang2017 ; Yin2018 . The driven dynamics during a linear quench across the localization transition point can be accurately described by the Kibble-Zurek scaling (KZS) Sinha2019 ; Liang2024 ; Xuan2023 .
In recent years, the novel phenomena induced by non-Hermiticity have been extensively studied Longhi2019 ; Luo2021 ; Jing2024 ; Yuto2021 ; Lv2022 ; Jiang2023 ; Han2023 ; Sarkar2023 ; Sunwy2024 ; XiaoL2024 ; Daitx2024 , including the non-Hermitian skin effect (NHSE) Zhang2109431 ; Liu2024 ; Gliozzi2024 ; YoshidaT2024 and exceptional points Wittrock2024 ; Wang024202 ; Martinez2018 ; Aodong706 ; Wingenbach2024 . As discussions on non-Hermitian mechanisms have deepened, researchers have integrated these mechanisms into the study of localization Jiang2019 ; Guo2021 ; Acharya2024 ; Guocx2024 ; Kochergin2024 . The introduction of non-Hermiticity can give rise to a new dimension for tuning localization transitions and uncovers a series of exotic new phenomena zhai2021 ; Jiang2019 ; zhai2020 ; Hatano1996 ; Hatano1998 ; Luo0153523 ; Chen144208 ; Wang024514 ; Longhi224206 ; Jazaeri2001 ; PWang2019 ; Tang2021 ; Chaohua123048 ; YCWang2023 ; KSuthar2022 ; Jiang2024 ; Hamazaki2019 . For instance, the localization transition in the non-Hermitian quasi-periodic system is accompanied by a transition from real to complex eigenvalues, as well as a topological phase transition in the eigenvalue spectrum Liu2021 ; Jiang2019 ; zhai2020 ; Hamazaki2019 . And the NHSE suggests that changes in boundary conditions can result in substantial alterations in the bulk localization properties of non-Hermitian systems Guo2021 . In addition, non-Hermiticity can also lead to critical exponents of the system belonging to different universality classes compared to the Hermitian case Luo2021 ; Zhai2022 .
Figure 1: (a) A schematic representation of the phase diagram for the non-Hermitian DAA model. When $\delta=-2J_{R}$ (denoted by the green point), this model recovers the non-Hermitian Anderson model. The light blue region denotes the localization region, the blue region denotes the critical region, the red point denotes the critical point of the non-Hermitian DAA model. The purple line depicting the skin effect phase under OBC or the extened phase under PBC. The typical spatial distributions of the ground eigenstates for the non-Hermitian DAA model are shown for $\Delta=0$ under (b1) OBC and (b2) PBC. Similarly, the typical spatial distributions of the ground eigenstates for $\Delta=1$ are illustrated under (c1) OBC and (c2) PBC. In these simulations, we use $g=0.5$, $\phi=0$, $\delta=-0.5$, and $L=377$.
Furthermore, the interplay between non-Hermiticity and non-equilibrium dynamics has also been studied Zhou2021 ; Xuan2023 ; Zhai2022 ; Zhaifphy ; Cheng2399 ; Xue2024 ; Zhang2023131 ; You2024109 ; Mak2024 ; Lizhen2024 . For example, researchers studied the dynamical evolutions of the non-Hermitian AA model with a complex incommensurate lattice. They found that Loschmidt echo dynamics cannot detect dynamical phase transitions when the post-quench parameter is in the PT symmetry-broken regime Huzh2021 . Another example is that by applying Floquet time-periodic driving fields to non-Hermitian quasicrystals, one can dynamically control localization transitions and mobility edges as the field parameters vary Zhou2021 . Besides, studies have shown that the KZS remains valid in describing the driven dynamics of the non-Hermitian AA model under both open boundary conditions (OBC) Zhaifphy and periodic boundary conditions (PBC) Zhai2022 .
More recently, researchers have increasingly focused on exploring the nonequilibrium dynamics of systems with multiple localization mechanisms Liang2024 ; Xuan2023 . Bu et al. introduced a Disorder-Aubry-André(DAA) model, combining disorder and quasi-periodic localization, and investigated its driven dynamics Xuan2022 ; Xuan2023 . Their work proposed a new scaling mechanism, providing a novel perspective on the study of localization phase transition dynamics in systems with coexisting multiple localization mechanisms. However, the interplay between the non-Hermiticity and the non-equilibrium dynamics in such systems remains unexplored. Here, we investigate the driven dynamics of localization transitions in a non-Hermitian DAA model with non-reciprocal hopping under both OBC and PBC. The KZS for the driven dynamics in the non-Hermitian DAA model is constructed and verified numerically across different quench directions. The rest of the paper is arranged as follows: the non-Hermitian DAA model and phase diagram are introduced in Sec. II. The static scaling properties are studied in Sec. III. Then, in Sec. IV, the KZS for the driven dynamics in the non-Hermitian DAA model under OBC and PBC is constructed and numerically verified. A summary is given in Sec. V.
II The non-Hermitian DAA Model and phase diagram
The Hamiltonian of the non-Hermitian DAA model is defined by Sun2024
$H$ $=$ $-\sum_{j}^{L}{(J_{L}c_{j}^{\dagger}c_{j+1}+J_{R}c_{j+1}^{\dagger}c_{j}})+\Delta\sum_{j}^{L}w_{j}c_{j}^{\dagger}c_{j}$
$+(2J_{R}+\delta)\sum_{j}^{L}\cos{[2\pi(\gamma j+\phi)]c_{j}^{\dagger}c_{j}}.$
Here, $c_{j}^{\dagger}(c_{j})$ are creation (annihilation) operator of the hard-core boson and $\gamma$ is an irrational number. $J_{L}=Je^{-g}$ and $J_{R}=Je^{g}$ are the asymmetry hopping coefficients. $\Delta$ and $(2J_{R}+\delta)$ measure the the amplitude of disorder and the quasi-periodic potential, respectively. $w_{j}\in[-1,1]$ gives the quenched disorder configuration, $\phi\in[0,1)$ is phase of the potential. We assume $J=1$ as the unity of energy, and $\gamma=(\sqrt{5}-1)/2$. To satisfy PBC of the quasi-periodic potential, we approximate $\gamma$ as a rational number $F_{n}/F_{n+1}$ where $F_{n+1}=L$ and $F_{n}$ are the Fibonacci numbers Jiang2019 ; Zhai2022 .
A schematic phase diagram of the non-Hermitian DAA model near the critical point is illustrated in Fig. 1(a). For $\delta=-2J_{R}$, this model returns to the non-Hermitian Anderson model Aubry1980 . However, when $\delta=0$, any finite $\Delta$ can also cause localization at the critical point of the non-Hermitian AA model, but the localization characteristics in this case differ from those of the purely non-Hermitian Anderson model. Therefore, for the non-Hermitian DAA model, the point $(\delta,\Delta)=(0,0)$ represents its critical point, and the critical region surrounding this point is jointly defined by $\Delta$ and $\delta$.
When $\Delta=0$, this model recovers the non-Hermitian AA model Jiang2019 . For $\delta>0$, all the eigenstates are localized under PBC and OBC. For $\delta<0$, all the eigenstates are extended under PBC Zhai2022 , whereas it exhibits the skin-effect phase under OBC Jiang2019 ; Zhaifphy . As illustrated in Figs. 1(b1) and (b2), the spatial distribution of the right eigenvector of the ground states ($|\Psi(j)\rangle$) with $\delta<0$ and $\Delta=0$ under OBC and PBC are depicted, respectively. Here, the ground state is defined as the eigenstate possessing the smallest real eigenenergy among all possible states. It is shown that the wave function is localized on the boundary under OBC, but it is evenly distributed within a small range under PBC. For $\delta>0$, the system will be in a localized state, regardless of the value of $\Delta$. For $\delta<0$, any finite $\Delta$ can also cause the system to transition into a localized state. As illustrated in Figs. 1(c1) and (c2), when $\Delta=1$ and $\delta<0$, the spatial distribution of localized states exhibits a high degree of similarity under both boundary conditions.
To discuss the impact of disorder on the localization phase transition in the non-Hermitian AA model, we further explore the energy spectrum of the non-Hermitian DAA model. Under PBC, the energy spectrum for the non-Hermitian AA model under PBC does not undergo significant changes when disorder is introduced, but it disrupts the correspondence between the transition from real to complex of energy spectrum and the localization transition Sun2024 . Here, Fig. 2 displays the energy spectrum of the non-Hermitian DAA model under OBC. The results show that, unlike the fully real energy spectrum of the non-Hermitian AA model under OBC Jiang2019 , the introduction of disorder leads to the emergence of imaginary parts in the energy spectrum in the region where $\delta<0$.
Figure 2: (a) Real and (b) imaginary parts of energy spectra of the model Eq.(1) under OBC. The black dashed line corresponds to $\delta=0$. Here, we choose $g=0.5$, $\phi=0.2$, $\Delta=0.8$ and $L=377$ in the calculation.
III Static critical properties of the Non-Hermitian DAA model
III.1 Static scaling forms in the critical region
In this section, we study the static properties within the critical region, utilizing the localization length ($\xi$), the inverse participation ratio ($\rm IPR$), and the energy gap between the ground state and the first excited state ($\Delta E$).
As in the usual quantum criticality, the localization length $\xi$ for the non-Hermitian system is given by Sinha2019 ; Zhai2022
$\xi=\sqrt{\sum_{n>n_{c}}^{L}[(n-n_{c})^{2}]P_{i}},$ (2)
where $P_{i}$ denotes the probability of the wavefunction at site $i$, and $n_{c}\equiv\sum nP_{i}$ represents the localization center. In the vicinity of a critical point, $\xi$ scales with the distance to the critical point $\varepsilon$ as Aubry1980
$\xi\propto\varepsilon^{-\nu},$ (3)
where $\nu$ is a critical exponent. For the pure non-Hermitian AA model, under both PBC and OBC, $\varepsilon=\delta$ and $\nu=\nu_{\delta}=1$ Zhai2022 ; Zhaifphy . For the pure non-Hermitian Anderson model, $\varepsilon=\Delta$ and $\nu=\nu_{A}=1.99$ Sun2024 . For the non-Hermitian DAA model, $\Delta$ constitutes another significant factor at $(\delta,\Delta)=(0,0)$ Xuan2022 ; Sun2024 . Specifically, when $\delta=0$ in Eq. (II), $\varepsilon=\Delta$ and $\nu=\nu_{\Delta}=0.52$ under PBC Sun2024 . $\nu_{\Delta}$ is a new parameter distinct from that of the non-Hermitian Anderson model and the non-Hermitian AA model.
The IPR is defined as Bauer1990 ; Fyodorov1992
${\rm IPR}=\frac{\sum_{j=1}^{L}||\Psi(j)\rangle|^{4}}{\sum_{j=1}^{L}||\Psi(j)\rangle|^{2}}.$ (4)
In the extended states, the scaling of ${\rm IPR}$ is given by ${\rm IPR}\propto L^{-1}$. For the localized states or skin effect sates, the scaling is ${\rm IPR}\propto L^{0}$ Jiang2019 . At the critical point, the $\rm IPR$ scales as
${\rm IPR}\propto L^{-s/\nu},$ (5)
where $s$ is a critical exponent. As $L$ approaches infinity, the ${\rm IPR}$ scales with the distance to the critical point $\varepsilon$ according to
${\rm IPR}\propto\varepsilon^{s}.$ (6)
For the non-Hermitian AA model, under both OBC and PBC, it is found that $\varepsilon=\delta$ and $s=s_{\delta}=0.1197$ Zhai2022 ; zhai2021 .
At the critical point of localization transition, the energy gap $\Delta E$ exhibits a scaling relationship with the lattice size $L$, given by
$\Delta E\propto L^{-z},$ (7)
where $z$ is a critical exponent. For the non-Hermitian AA model, under both OBC and PBC, it is found that $\varepsilon=\delta$ and $z=z_{\delta}=2$ Zhai2022 . when $L$ tends to infinity, energy gap $\Delta E$ scales as
$\Delta E\propto\varepsilon^{\nu z}.$ (8)
Considering the finite-size effect, the general scaling formulation for a quantity $Y$ referring to a single parameter can be expressed as:
$Y(\varepsilon)$ $=$ $L^{y/\nu}f(\varepsilon L^{1/\nu}),$ (9)
where $y$ is the critical exponent of $Y$, defined such that $Y\propto\varepsilon^{-y}$ as $L\rightarrow\infty$. The function $f(.)$ represents the scaling function. By taking $\Delta$ and $\delta$ as scaling variables simultaneously, the comprehensive finite-size scaling expressions for the three observables are given by
$\xi$ $=$ $Lf_{1}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}),$ (10)
${\rm IPR}$ $=$ $L^{-s_{\delta}/\nu_{\delta}}f_{2}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}),$ (11)
${\Delta E}$ $=$ $L^{-z_{\delta}}f_{3}(\delta L^{1/\nu_{\delta}},\Delta L^{1/\nu_{\Delta}}).$ (12)
These scaling functions, Eqs. (10) to (12), are applicable within the critical region of the non-Hermitian DAA model.
Figure 3: Scaling properties in the ground state for fixed $\delta L^{1/\nu_{\delta}}=1$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for $1000$ samples. The double-logarithmic scales are used.
III.2 Numerical results of static scaling of the non-Hermitian DAA model
The critical exponents of the non-Hermitian DAA model under PBC have been examined Sun2024 . It is found that the critical exponents vary along the $\delta$ and $\Delta$ directions because these parameters represent two distinct relevant directions. Under PBC, the critical exponents along $\delta$ direction with $\Delta=0$ are $(\nu,s,z)=(\nu_{\delta},s_{\delta},z_{\delta})=(1,0.1197,2)$ Zhai2022 . In contrast, the scaling exponents along the $\Delta$ direction are $(\nu,s,z)=(\nu_{\Delta},s_{\Delta},z_{\Delta})=(0.52,0.0642,2)$.
Figure 4: Scaling properties in the ground state for fixed $\delta L^{1/\nu_{\delta}}=-1$ under OBC. The curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaled for different $L^{\prime}$s. The curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaled for different $L^{\prime}$s. The curves of $\Delta E$ versus $\Delta$ before (c1) and after (c2) rescaled for different $L^{\prime}$s. Here, $g=0.5$, and the result is averaged for $1000$ samples. The double-logarithmic scales are used.
Here, we study the static scaling properties of the non-Hermitian DAA model under OBC by employing the exponents obtained from fitting under PBC. We numerically validate Eqs. (10) to (12) under OBC by maintaining $\delta L^{1/\nu_{\delta}}$ at a constant value. Fig. 3 depict the scaling behavior of $\xi$, $\rm IPR$ and $\Delta E$ as functions of $\Delta$ for $\delta L^{1/\nu_{\delta}}=1$. After rescaling these values according to Eqs. (10) to (12), the resulting curves align closely, thereby confirming the validity of these equations. For the critical region where $\delta<0$, similar numerical results are presented in Fig. 4 for $\delta L^{1/\nu_{\delta}}=-1$. The alignment of the rescaled curves shown in Figs. 4 (a2), (b2) and (c2) further verifies Eqs. (10) to (12).
These results confirm that the same set critical exponent of the non-Hermitian DAA model under PBC is typically sufficient to characterize the critical behavior in the critical region under OBC.
IV KZS for the driven dynamics of the non-Hermitian DAA model
Figure 5: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
IV.1 General theory of the KZS
We proceed to investigate the KZS associated with the driven dynamics within the non-Hermitian DAA model. Our analysis assumes that the system begins in a localized state and is subsequently driven through the critical region by linearly adjusting the distance $\varepsilon$ over time $t$ with a constant rate $R$. The temporal evolution of $\varepsilon$ is described by
$\varepsilon=\varepsilon_{0}-Rt,$ (13)
where $\varepsilon$ can represent either $\Delta$ or $\delta$, depending on the specific situation being analyzed, $\varepsilon_{0}>0$ denotes the initial distance from the critical point at $t=0$. According to the KZS, if $|\varepsilon|>R^{1/r\nu}$ with the scaling exponent $r=z+1/\nu$, the system has enough time to adapt to changes in the Hamiltonian, maintaining adiabatic conditions. Conversely, when $|\varepsilon|<R^{1/r\nu}$, the system’s internal change rate lags behind the external parameter’s rate, indicating the system entry into the impulse region Zhong1 ; Zhong2 ; Zhong3 . Below, we choose the system size $L=610$, which is sufficiently large to disregard finite-size effects in realtime simulations. We initialize the system in its ground state with a fixed $\varepsilon_{0}$ value of 1.0.
Figure 6: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta=0$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
Apparently, there are two adjustable variables $\Delta$ and $\delta$ in the non-Hermitian DAA model. Thus, both of these relevant terms should be incorporated into the full KZS form. The evolution of the localization length $\xi$ should satisfy Xuan2023 ; Liang2024
$\xi(\Delta,\delta,R)=R^{-\frac{1}{r_{\Delta}}}f_{4}(\delta R^{-\frac{1}{r_{\Delta}\nu_{\delta}}},\Delta R^{-\frac{1}{r_{\Delta}\nu_{\Delta}}}),$ (14)
where $r_{\Delta}=z_{\Delta}+1/\nu_{\Delta}$. Similarly, $\rm IPR$ should satisfy
${\rm IPR}(\Delta,\delta,R)=R^{\frac{s_{\delta}}{r_{\Delta}\nu_{\delta}}}f_{5}(\delta R^{-\frac{1}{r_{\Delta}\nu_{\delta}}},\Delta R^{-\frac{1}{r_{\Delta}\nu_{\Delta}}}).$ (15)
Eq. (14) and Eq. (15) should be applicable for a range of $\Delta$ and $\delta$ values in proximity to the critical region of the non-Hermitian DAA model.
In order to verify the KZS, we can either set some variables to zero or fix the values of certain variables. When $\Delta=0$, the model (II) corresponds to the non-Hermitian AA model, and the KZS for the non-Hermitian AA model have been validated Zhai2022 ; Zhaifphy . When $\delta=0$, Eq. (14) and Eq. (15) return to
$\xi(\Delta,R)$ $=$ $R^{-1/r_{\Delta}}f_{6}(\Delta R^{-1/r_{\Delta}\nu_{\Delta}}),$ (16)
${\rm IPR}(\Delta,R)$ $=$ $R^{s_{\delta}/r_{\Delta}\nu_{\delta}}f_{7}(\Delta R^{-1/r_{\Delta}\nu_{\Delta}}).$ (17)
This corresponds to the situation where a dynamic quench of the system is conducted along the direction of $\Delta$, with $\delta$ fixed at 0. For the case where $\delta\neq 0$, we can also fix $\delta R^{-1/r_{\Delta}\nu_{\delta}}$, and $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ to verify Eq. (14) and Eq. (15).
Figure 7: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under OBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
IV.2 Numerical results of driven dynamics
First, we verify the scaling function Eq. (16) and Eq. (17) with the initial state deep in the localization state by linearly varying the value of the disorder potential $\Delta$. The numerical results for $\xi$ and $\rm IPR$ as functions of $\Delta$ for various driving rates $R$ under OBC are depicted in Fig. 5(a1) and Fig. 5(b1), respectively. Observation reveals that the curves corresponding to different $R$ diverge from each other near the critical region. However, when these curves are rescaled using the critical exponents of the non-Hermitian DAA model, they converge near the critical region, as illustrated in Fig. 5(a2) and Fig. 5(b2). This convergence aligns with the predictions of Eq. (16) and Eq. (17). Additionally, in Fig. 6, the dynamics of $\xi$ and $\rm IPR$ under PBC are plotted. Similarly, the rescaled curves collapse onto each other, confirming Eq. (16) and Eq. (17).
Then, we verify the KZS forms of $\xi$ and $\rm IPR$ by fixing $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$, which is demonstrated in Fig. 7. We numerically compute the time evolutions of $\xi$ and $\rm IPR$ for various driving rate $R$ under OBC, with results presented in Fig. 7(a1) and Fig. 7(b1), respectively. Upon rescaling the quantities according to Eq. (14) and Eq. (15), we observe that the curves corresponding to different $R$ converge near the critical region in Fig. 7(a2) and Fig. 7(b2). Similar results for $\xi$ and $\rm IPR$ under PBC are plotted in Fig. 8. After rescaling with $R$, the rescaled curves align with each other as shown in Fig. 8(a2) and Fig. 8(b2), confirming Eq. (14) and Eq. (15).
Figure 8: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\delta R^{-1/r_{\Delta}\nu_{\delta}}=0.3$ under PBC. Curves of $\xi$ versus $\Delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\Delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
Figure 9: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under OBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
We have also numerically verified the KZS for the non-Hermitian DAA model by linearly varying the value of the quasiperiodic potential $\delta$. By fixing $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$, we calculate the evolution of $\xi$ and $\rm IPR$ for various driving rates $R$, starting from the initial state. By rescaling these evolution curves with respect to $R$, we observe that the rescaled curves align with each other, as illustrated in Fig. 9, confirming Eq. (14) and Eq. (15). In Fig. 10, we present the evolution of $\xi$ and $\rm IPR$ under PBC. After rescaling the curves converge onto one another, in accordance with Eq. (14) and Eq. (15). These findings demonstrate the validity of the KZS in both $\Delta$ and $\delta$ directions near the critical region of the non-Hermitian DAA model.
Furthermore, we also observe that for each of the three quenching paths we considered, there is basically no difference in the results of the dynamic numerical simulations between OBC and PBC. This phenomenon can be understood from the following two aspects. Firstly, all three dynamic evolution processes involve transitions from a localized state, passing through a critical region, and ending up in another localized state. The localized states are not sensitive to changes in boundary conditions between OBC and PBC. Secondly, for each of the three paths we considered, their energy spectra exhibit similar characteristics under both OBC and PBC. For the case where $\delta=0$ and $\Delta$ is varied, the energy spectra under both OBC and PBC are basically no difference; for the case where $\delta R^{-1/r_{\Delta}\nu_{\delta}}$ is fixed and $\Delta$ is varied, the energy spectra under both OBC and PBC are real; for the case where $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}$ is fixed and $\delta$ is varied, the energy spectrum undergoes a transition from real to imaginary under both OBC and PBC.
Figure 10: KZS of driven dynamics in the non-Hermitian DAA model with fixed $\Delta R^{-1/r_{\Delta}\nu_{\Delta}}=0.3$ under PBC. Curves of $\xi$ versus $\delta$ before (a1) and after (a2) rescaling for different driving rates $R$. Curves of $\rm IPR$ versus $\delta$ before (b1) and after (b2) rescaling for different $R$. The lattice size is $L=610$, $\phi=0$, $g=0.5$ and one sample of $w_{j}$ is used.
V summary
In summary, we have conducted a comprehensive study of the static scaling behavior and the driven dynamics of localization transitions in the non-Hermitian DAA model, considering both OBC and PBC. By examining the static behavior of $\xi$, $\rm IPR$ and $\Delta E$, we have demonstrated that the same critical exponents observed under PBC are also valid under OBC. We constructed the KZS for driven dynamics of the non-Hermitian DAA systems and numerically confirmed its validity across three different quench directions. For the dynamical paths we considered, the boundary conditions do not have a significant impact on the dynamical evolution process. Our work generalizes the application of the KZS to the dynamical localization transitions within systems featuring dual localization mechanisms.
On-site gain and loss constitute another crucial aspect of non-Hermiticity Tomasi2023 ; Longhi2023 , with numerous studies highlighting its distinct influence on localization compared to nonreciprocal effects. Hence, investigating the non-Hermitian DAA model that incorporates on-site gain and loss stands as a potential and fruitful extension of this paper. Additionally, the absence of a mobility edge in the excited states of DAA model has been demonstrated Xuan2022 ; Sun2024 . However, the introduction of dynamic driving to systems that do possess a mobility edge may result in unique phenomena, necessitating further exploration and discussion.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 12274184), the Qing Lan Project, the National Natural Science Foundation of China (Grant No. 12404105) and the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (Grant No. 24KJB140008).